# Test-time regression: a unifying framework for designing sequence models with associative memory

### v1

@@ Abstract
We first explain the mathematical notation we will use in our derivations. We use bold lowercase letters to refer to vectors (e.g. $\mathbf{x},\mathbf{k},\mathbf{v}$) and bold uppercase letters to refer to matrices (e.g. $\mathbf{M},\mathbf{K},\mathbf{V}$). Assume we have a sequence of $T$ inputs $\mathbf{x}_{1},\ldots,\mathbf{x}_{T}$ which have already transformed into a sequence of keys and values $(\mathbf{k}_{1},\mathbf{v}_{1}),\ldots,(\mathbf{k}_{T},\mathbf{v}_{T})$. We will use these keys and values as the dataset for our test-time regression problem. Let $\mathbf{K}_{t}=\begin{bmatrix}\mathbf{k}_{1}\ldots\mathbf{k}_{t}\end{bmatrix}{}^{\top}\in\mathbb{R}^{t\times{D_{k}}}$ and $\mathbf{V}_{t}=\begin{bmatrix}\mathbf{v}_{1}\ldots\mathbf{v}_{t}\end{bmatrix}{}^{\top}\in\mathbb{R}^{t\times{D_{v}}}$ be matrices that contain $t$ keys and $t$ values respectively, for $t=1,\ldots,T$. Each key/value vector is a row of the key/value matrix. To minimize clutter, we will use the shorthand $\mathbf{K}$ and $\mathbf{V}$ instead of $\mathbf{K}_{t}$ and $\mathbf{V}_{t}$ when the timestep $t$ is clear from context.
@@
    Choice of weights: assign equal weight to each association,
@@
$\mathbf{M}_{t}=\operatorname*{\arg\!\min}_{\mathbf{M}}\frac{1}{2}\sum_{i=1}^{t}\lVert\mathbf{v}_{i}-\mathbf{M}\mathbf{k}_{i}\rVert_{2}^{2}=\begin{cases}\mathbf{V}{}^{\top}\mathbf{K}(\mathbf{K}{}^{\top}\mathbf{K})^{-1},\ \text{overdetermined, e.g. }t\geq{D_{k}},\\ \mathbf{V}{}^{\top}(\mathbf{K}\mathbf{K}{}^{\top})^{-1}\mathbf{K},\ \text{underdetermined, e.g. }t<{D_{k}},\end{cases}$ (3)
where we choose the min-norm solution in the underdetermined case. The overdetermined case corresponds to $\mathbf{K}$ having linearly-independent columns, which is not possible if $t<{D_{k}}$, while the underdetermined case corresponds to $\mathbf{K}$ having linearly-independent rows, which is not possible if $t>{D_{k}}$. We restrict our discussion to the case when $t\geq{D_{k}}$, which is the case for typical applications. Applying Equation 3 as a layer was previously explored by von Oswald et al. (2023) (who called it a mesa-layer) in the context of understanding deep transformers and by Garnelo and Czarnecki (2023) (who called an intention layer) on the expressivity of self-attention, but here we derive it as mathematical consequence of requiring the sequence model to be able to perform associative recall.
@@
Although ${m}_{\textsc{linear},\,t}$ is the optimal linear memory with respect to our objective function, it can be expensive to compute, requiring us to invert a ${D_{k}}\times{D_{k}}$ matrix at every timestep. One crude but hardware-efficient simplification would be to assume that $\mathbf{K}{}^{\top}\mathbf{K}=\mathbf{I}$, avoiding matrix inversion completely, at the cost of some recall ability. After making this rough approximation, we arrive exactly at the equations for linear attention: $\mathbf{y}_{t}=\mathbf{M}_{t}\mathbf{q}_{t}\approx\mathbf{V}{}^{\top}\mathbf{K}\mathbf{q}_{t}=\sum_{i=1}^{t}\mathbf{v}_{i}\mathbf{k}_{i}{}^{\top}\mathbf{q}_{t}$.
Our derivation clarifies exactly how linear attention fails as an associative memory. When the keys are zero mean, $\mathbf{K}{}^{\top}\mathbf{K}$ is proportional to the empirical covariance matrix of the keys. Thus linear attention is a crude associative memory that ignores the covariance between the dimensions of the key vectors. In fact, it is an optimal linear memory only when $t\leq{D_{k}}$ and the keys are orthonormal $\mathbf{k}_{i}{}^{\top}\mathbf{k}_{j}=\delta_{ij}$.
From a training stability perspective, the inverse covariance factor helps prevent the output $\mathbf{y}_{t}$ from becoming unbounded. We can bound the norm of the output of a least squares layer (Equation 3) by $\lVert\mathbf{y}_{t}\rVert_{2}=\lVert\mathbf{V}{}^{\top}\mathbf{K}(\mathbf{K}{}^{\top}\mathbf{K})^{-1}\mathbf{q}_{t}\rVert_{2}\leq\lVert\mathbf{q}_{t}\rVert_{2}\left(\sum_{i=1}^{t}\lVert\mathbf{v}_{i}\rVert_{2}\lVert\mathbf{k}_{i}\rVert_{2}\right)/\lambda_{\text{min}}(\mathbf{K}{}^{\top}\mathbf{K})$ where $\lambda_{\text{min}}(\mathbf{K}{}^{\top}\mathbf{K})$ is the minimum eigenvalue of $\mathbf{K}{}^{\top}\mathbf{K}$. See Appendix A for the derivation. When we ignore the covariance matrix, approximating it with the identity, the denominator disappears and the output $\mathbf{y}_{t}$ can grow arbitrarily large with the sequence length, leading to training instability. This instability indeed shows up in practice for linear attention, and Qin et al. (2022) proposed output normalization to stabilize training. Our derivation shows that output normalization works by approximately restoring the self-normalizing property of Equation 3 that linear attention loses. We will see in Section 4 that ignoring the covariance causes linear attention to perform worse at associative recall.
@@
Conveniently, we can compute $\mathbf{M}_{t}$ in Equation 3 for $t=1,\ldots T$ in parallel or sequentially. To compute them in parallel, we solve a batch of $T$ independent linear systems. To compute them sequentially, we note that $(\mathbf{K}_{t}{}^{\top}\mathbf{K}_{t})^{-1}=(\mathbf{K}_{t-1}{}^{\top}\mathbf{K}_{t}+\mathbf{k}_{t}\mathbf{k}_{t}{}^{\top})^{-1}$, allowing us to autoregressively compute the inverses via the Woodbury matrix identity. The sequential form is known as online linear regression or recursive least squares (RLS) (Haykin, 2014). Thus during training we can use the parallel form, and during testing we can unroll the recurrence into a recurrent neural network. In prior deep learning literature, this kind of unweighted linear regression layer has been called an intention layer (Garnelo and Czarnecki, 2023) and a mesa-layer (von Oswald et al., 2023), though not motivated by associative memory.
@@
So far we have only considered linear associative memory implemented by linear functions ${\mathcal{M}_{\textsc{linear}}}$ which have limited expressiveness. We can derive nonlinear memories by using a feature map $\phi:\mathbb{R}^{D_{k}}\to\mathbb{R}^{D_{\phi}}$, replacing each key $\mathbf{k}_{i}$ with $\phi(\mathbf{k}_{i})$. Then our function class is defined by $\mathcal{M}_{\phi}=\{{m}\mid{m}(\mathbf{k})=\mathbf{M}\phi(\mathbf{k}),\ \mathbf{M}\in\mathbb{R}^{{D_{v}}\times{D_{\phi}}}\}$ and we solve the least-squares objective $\operatorname*{\arg\!\min}_{\mathbf{M}}\sum_{i=1}^{t}\lVert\mathbf{v}_{i}-\mathbf{M}\phi(\mathbf{k}_{i})\rVert_{2}^{2}/2$. Solving this least squares objective for $\mathbf{M}$ results in an optimal associative memory that is nonlinear with respect to the input keys. Parameterizing associative memory with a nonlinear feature map has a long history, dating back to Poggio (1975) which found polynomial feature maps to be highly effective, consistent with more recent works (Zhang et al., 2023, Arora et al., 2024).
If we make the same crude approximation as linear attention of disregarding the covariance among $\{\phi(\mathbf{k}_{i})\}_{i=1}^{t}$, we arrive at a suboptimal nonlinear memory. Thus, recent works that apply linear attention with a nonlinear feature map, such as Schlag et al. (2021), Kasai et al. (2021), Zhang et al. (2023), Chen et al. (2024), and others, can be understood within our framework as suboptimal nonlinear memory as well.
@@
Classic regression theory tells us that unweighted linear regression works best when the data is stationary (Hastie et al., 2009, Haykin, 2014). However, time-series data and text data are typically non-stationary, and recent tokens are more informative of predicting the next token. Thus a natural requirement is to encourage our memory ${m}_{t}$ to capture more of the recent history by downweighting the past. Fortunately, this can be straightforwardly enforced through classic weighted linear least squares regression together with geometrically decaying weights to forget older key-value pairs.
To develop an architecture from weighted linear least squares, we change only the optimization objective relative to regular linear least squares:
@@
    where the weights are geometrically decaying in time: $\gamma_{i}^{(t)}=\prod_{j=i+1}^{t}\gamma_{j}$ and $\gamma_{i}\in[0,1]$ for $i=1,\ldots,t$.
@@
Here the analytical min-norm solution is
$\mathbf{M}_{t}$ =$=\begin{cases}\mathbf{V}{}^{\top}\bm{\Gamma}\mathbf{K}(\mathbf{K}{}^{\top}\bm{\Gamma}\mathbf{K})^{-1},\ \text{overdetermined, e.g. }\,T\geq{D_{k}}\\ \mathbf{V}{}^{\top}(\mathbf{K}\mathbf{K}{}^{\top})^{-1}\mathbf{K},\ \text{underdetermined, e.g. }\ T<{D_{k}}.\end{cases}$ (5)
where we defined a diagonal matrix $\bm{\Gamma}$ (omitting the $t$ subscript following our convention) with entries $\bm{\Gamma}_{ii}=\gamma_{i}^{(t)}$. The underdetermined case has the same solution in both the weighted and unweighted case.
@@
Once again, computing $\mathbf{M}_{t}$ requires matrix inversion which is not efficient on matrix-multiplication based hardware. Making the same crude approximation as before of $\mathbf{K}{}^{\top}\bm{\Gamma}\mathbf{K}\approx\mathbf{I}$ results in $\mathbf{M}_{t}\approx\mathbf{V}{}^{\top}\bm{\Gamma}\mathbf{K}=\sum_{i=1}^{t}\gamma_{i}^{(t)}\mathbf{v}_{i}\mathbf{k}_{i}{}^{\top}$. We can further unroll this equation into a recurrence due to the multiplicative structure of our weights:
@@
This is exactly the recurrence implemented by gated variants of linear attention, such as GLA (Yang et al., 2024b), RetNet (Sun et al., 2023), Gateloop (Katsch, 2024), RWKV-6 (Peng et al., 2024), mLSTM (Beck et al., 2024), LRU (Orvieto et al., 2023), as well as state-space models like Mamba-2 (Dao and Gu, 2024). Forgetting older tokens significantly improves performance, which was first discovered empirically (Hochreiter and Schmidhuber, 1997, Greff et al., 2017, van der Westhuizen and Lasenby, 2018). Our derivation of gating from weighted least squares provides a mathematically grounded explanation for the benefits of forgetting through the classic theory of online regression on non-stationary data.
@@
A computationally cheaper alternative is to apply a first-order gradient-descent method, corresponding to the following design choices:
@@
    Choice of weights: assign equal weight to each association,
@@
$\nabla\mathcal{L}_{t}(\mathbf{M})$ $=\sum_{i=1}^{t}\nabla L_{i}(\mathbf{M})=\sum_{i=1}^{t}(\mathbf{M}\mathbf{k}_{i}-\mathbf{v}_{i})\mathbf{k}_{i}{}^{\top}=(\mathbf{M}\mathbf{K}{}^{\top}-\mathbf{V}{}^{\top})\mathbf{K}.$ (9)
to find $\mathbf{M}_{t}$. Let the initialization be at the origin $\mathbf{M}^{(0)}_{t}=\mathbf{0}$. Iteratively solving this objective with full-batch gradient descent with step size $\beta_{i}$ at each iteration results in
$\mathbf{M}^{(i)}_{t}$ $=\mathbf{M}^{(i-1)}_{t}-\beta_{i}\nabla\mathcal{L}_{t}(\mathbf{M}^{(i-1)}_{t})$ (10)
$=\mathbf{M}^{(i-1)}_{t}-\beta_{i}\sum_{j=1}^{t}\nabla L_{j}(\mathbf{M}^{(i-1)}_{t})$ (11)
$=\mathbf{M}^{(i-1)}_{t}-\beta_{i}\sum_{j=1}^{t}(\mathbf{M}^{(i-1)}_{t}\mathbf{k}_{j}-\mathbf{v}_{j})\mathbf{k}_{j}{}^{\top}$ (12)
$=\mathbf{M}^{(i-1)}_{t}(\mathbf{I}-\sum_{j=1}^{t}\beta_{i}\mathbf{k}_{j}\mathbf{k}_{j}{}^{\top})+\beta_{i}\sum_{j=1}^{t}\mathbf{v}_{j}\mathbf{k}_{j}{}^{\top}.$ (13)
@@
Preconditioning gradient descent.
Consider the $I=1$ case with step size $\beta_{i}=1$. Since we initialize at the origin, we obtain the equation for linear attention
$\mathbf{M}^{(1)}_{t}=\mathbf{M}^{(0)}_{t}-\nabla\mathcal{L}_{t}(\mathbf{M}^{(0)}_{t})=\mathbf{V}{}^{\top}\mathbf{K},$ (14)
a result also shown by (Sun et al., 2024, Irie et al., 2022).
A standard technique to accelerate the convergence of gradient descent is to precondition the gradient (Boyd and Vandenberghe, 2004). The optimal preconditioner in this case, due to the quadratic objective, is the Jacobian of $\mathcal{L}_{t}$, $\mathbf{J}_{\mathcal{L}_{t}}(\mathbf{M})=\mathbf{K}{}^{\top}\mathbf{K}$. This preconditioner allows gradient descent to converge in a single update
$\mathbf{M}^{(1)}_{t}=\mathbf{M}^{(0)}_{t}-\nabla\mathcal{L}_{t}(\mathbf{M}^{(0)}_{t})\mathbf{J}_{\mathcal{L}_{t}}(\mathbf{M}^{(0)}_{t})^{-1}=\mathbf{V}{}^{\top}\mathbf{K}(\mathbf{K}{}^{\top}\mathbf{K})^{-1},$ (15)
@@
By taking this iterative optimization perspective, we see that linear attention and Equation 3 lie on two extremes: batch gradient descent with no preconditioning versus batch gradient descent with perfect preconditioning. The difference in performance comes from whether the optimizer accounts for the curvature of the objective, governed by the covariance of the keys (Boyd and Vandenberghe, 2004). Exploring other preconditioners, e.g. quasi-Newton methods, that interpolates between the two extremes may be an interesting direction for future research.
Online learners as stochastic gradient descent.
@@
An alternative to full-batch gradient descent is stochastic gradient descent (SGD) where we use a subset of the samples to compute the gradient at each step. This is particularly convenient in causal sequence modeling where keys and values typically are presented one at a time, rather than all at once. In this streaming setting, we can instead perform single-example SGD to iteratively minimize Equation 1. In the case of linear memories ${\mathcal{M}_{\textsc{linear}}}$, Equation 11 with a single-example becomes
$\mathbf{M}_{t}$ $=\mathbf{M}_{t-1}-\beta_{t}\nabla L_{t}(\mathbf{M}_{t-1})$ (16)
$=\mathbf{M}_{t-1}+\beta_{t}(\mathbf{v}_{t}-\mathbf{M}_{t-1}\mathbf{k}_{t})\mathbf{k}_{t}{}^{\top}$ (17)
$=\mathbf{M}_{t-1}(\mathbf{I}-\beta_{t}\mathbf{k}_{t}\mathbf{k}_{t}{}^{\top})+\beta_{t}\mathbf{v}_{t}\mathbf{k}_{t}{}^{\top},$ (18)

### v3

@@ Abstract
We first explain the mathematical notation we will use in our derivations. We use bold lowercase letters to refer to vectors (e.g. $\mathbf{x},\mathbf{k},\mathbf{v}$) and bold uppercase letters to refer to matrices (e.g. $\mathbf{M},\mathbf{K},\mathbf{V}$). Assume we have a sequence of $T$ inputs $\mathbf{x}_{1},\ldots,\mathbf{x}_{T}$ which have already transformed into a sequence of keys and values $(\mathbf{k}_{1},\mathbf{v}_{1}),\ldots,(\mathbf{k}_{T},\mathbf{v}_{T})$. We will use these keys and values as the dataset for our test-time regression problem. Let $\mathbf{K}_{t}=\begin{bmatrix}\mathbf{k}_{1}\ldots\mathbf{k}_{t}\end{bmatrix}{}^{\top}\in\mathbb{R}^{t\times{D_{k}}}$ and $\mathbf{V}_{t}=\begin{bmatrix}\mathbf{v}_{1}\ldots\mathbf{v}_{t}\end{bmatrix}{}^{\top}\in\mathbb{R}^{t\times{D_{v}}}$ be matrices that contain $t$ keys and $t$ values respectively, for $t=1,\ldots,T$. Each key/value vector is a row of the key/value matrix.
@@
    Choice of weights: assign equal weight $\gamma_{i}^{(t)}=1$ to each association,
@@
$\mathbf{M}_{t}=\operatorname*{\arg\!\min}_{\mathbf{M}}\frac{1}{2}\sum_{i=1}^{t}\lVert\mathbf{v}_{i}-\mathbf{M}\mathbf{k}_{i}\rVert_{2}^{2}=\begin{cases}\mathbf{V}_{t}{}^{\top}\mathbf{K}_{t}(\mathbf{K}_{t}{}^{\top}\mathbf{K}_{t})^{-1},\ \text{overdetermined, e.g. }t\geq{D_{k}},\\ \mathbf{V}_{t}{}^{\top}(\mathbf{K}_{t}\mathbf{K}_{t}{}^{\top})^{-1}\mathbf{K}_{t},\ \text{underdetermined, e.g. }t<{D_{k}},\end{cases}$ (11)
where we choose the min-norm solution in the underdetermined case. The overdetermined case corresponds to $\mathbf{K}_{t}$ having rank ${D_{k}}$ (linearly independent columns) and the underdetermined case corresponds to $\mathbf{K}_{t}$ having rank $t$ (linearly independent rows). In general, we can express the solution in terms of a Moore-Penrose pseudoinverse: $\mathbf{M}_{t}=\mathbf{V}_{t}{}^{\top}(\mathbf{K}_{t}^{\dagger}){}^{\top}$. We restrict our discussion to the case of $t\geq{D_{k}}$ with linearly independent columns, which is the case for typical applications.
@@
Although ${m}_{\textsc{linear},\,t}$ is the optimal linear memory with respect to our objective function, it can be expensive to compute, requiring us to invert a ${D_{k}}\times{D_{k}}$ matrix at every timestep. One crude but hardware-efficient simplification would be to assume that $\mathbf{K}_{t}{}^{\top}\mathbf{K}_{t}=\mathbf{I}$, avoiding matrix inversion completely, at the cost of some recall ability. After making this rough approximation, we arrive exactly at the equations for linear attention:
$\mathbf{y}_{t}={m}_{\textsc{linear}}(\mathbf{q}_{t})=\mathbf{V}_{t}{}^{\top}\mathbf{K}_{t}(\mathbf{K}_{t}{}^{\top}\mathbf{K}_{t})^{-1}\mathbf{q}_{t}\approx\mathbf{V}_{t}{}^{\top}\mathbf{K}_{t}\mathbf{q}_{t}=\sum_{i=1}^{t}\mathbf{v}_{i}\mathbf{k}_{i}{}^{\top}\mathbf{q}_{t}.$ (12)
Our derivation clarifies exactly how linear attention underperforms as an associative memory. When the keys are zero mean, $\mathbf{K}_{t}{}^{\top}\mathbf{K}_{t}$ is proportional to the empirical covariance matrix of the keys. Thus linear attention is a crude associative memory that ignores the covariance between the dimensions of the key vectors. In fact, it is an optimal linear memory only when $t\leq{D_{k}}$ and the keys are orthonormal $\mathbf{k}_{i}{}^{\top}\mathbf{k}_{j}=\delta_{ij}$.
From a training stability perspective, the inverse covariance factor helps prevent the output $\mathbf{y}_{t}$ from becoming unbounded. We can bound the norm of the output of a least squares layer (Equation 11) by
$\lVert\mathbf{y}_{t}\rVert_{2}=\lVert\mathbf{V}_{t}{}^{\top}\mathbf{K}_{t}(\mathbf{K}_{t}{}^{\top}\mathbf{K}_{t})^{-1}\mathbf{q}_{t}\rVert_{2}\leq\frac{\lVert\mathbf{q}_{t}\rVert_{2}\left(\sum_{i=1}^{t}\lVert\mathbf{v}_{i}\rVert_{2}\lVert\mathbf{k}_{i}\rVert_{2}\right)}{\lambda_{\text{min}}(\mathbf{K}_{t}{}^{\top}\mathbf{K}_{t})}$ (13)
where $\lambda_{\text{min}}(\mathbf{K}_{t}{}^{\top}\mathbf{K}_{t})$ is the minimum eigenvalue of $\mathbf{K}_{t}{}^{\top}\mathbf{K}_{t}$. See Appendix A for the derivation. When we ignore the covariance matrix, approximating it with the identity, the denominator disappears and the output $\mathbf{y}_{t}$ can grow arbitrarily large with the sequence length, leading to training instability. This instability indeed shows up in practice for linear attention, and Qin et al. (2022) proposed output normalization to stabilize training. Our derivation shows that output normalization works by approximately restoring the self-normalizing property of Equation 11 that linear attention loses.
Regardless of training stability, we will see in Section 6 that ignoring the covariance causes linear attention to perform worse at associative recall.
@@
Conveniently, we can compute $\mathbf{M}_{t}$ in Equation 11 for $t=1,\ldots T$ in parallel or sequentially. To compute them in parallel, we solve a batch of $T$ independent linear systems. To compute them sequentially, we note that $(\mathbf{K}_{t}{}^{\top}\mathbf{K}_{t})^{-1}=(\mathbf{K}_{t-1}{}^{\top}\mathbf{K}_{t}+\mathbf{k}_{t}\mathbf{k}_{t}{}^{\top})^{-1}$, allowing us to autoregressively compute the inverses via the Woodbury matrix identity. The sequential form is known as online linear regression or recursive least squares (RLS) (Haykin, 2014). Thus during training we can parallelize across the sequence axis by solving a batch of $T$ linear systems, and during testing we can unroll the recurrence into a recurrent neural network. In prior deep learning literature, this idea of linear-regression-as-a-layer has been explored by Garnelo and Czarnecki (2023) and von Oswald et al. (2023) under the names “intention” and “mesa-layer”, respectively, motivated by heuristic similarities to softmax attention, rather than deriving them from first principles via enforcing associative recall as we do here.
@@
So far we have only considered linear associative memory implemented by linear functions in ${\mathcal{M}_{\textsc{linear}}}$, which have limited expressiveness. We can derive nonlinear memory by using a feature map $\phi:\mathbb{R}^{D_{k}}\to\mathbb{R}^{D_{\phi}}$, replacing each key $\mathbf{k}_{i}$ with $\phi(\mathbf{k}_{i})$. Then our function class is defined by $\mathcal{M}_{\phi}=\{{m}\mid{m}(\mathbf{k})=\mathbf{M}\phi(\mathbf{k}),\ \mathbf{M}\in\mathbb{R}^{{D_{v}}\times{D_{\phi}}}\}$ and we solve the least-squares objective $\operatorname*{\arg\!\min}_{\mathbf{M}}\sum_{i=1}^{t}\lVert\mathbf{v}_{i}-\mathbf{M}\phi(\mathbf{k}_{i})\rVert_{2}^{2}/2$. Solving this least squares objective for $\mathbf{M}$ results in an optimal associative memory that is nonlinear with respect to the input keys. Parameterizing associative memory with a nonlinear feature map has a long history, dating back to Poggio (1975) which found polynomial feature maps to be highly effective, consistent with more recent works (Zhang et al., 2023a, Arora et al., 2024).
If we make the same crude approximation as linear attention of disregarding the covariance among $\{\phi(\mathbf{k}_{i})\}_{i=1}^{t}$, we arrive at a suboptimal nonlinear memory. Thus, recent works that apply linear attention with a nonlinear feature map, such as Schlag et al. (2021), Kasai et al. (2021), Zhang et al. (2023a), Chen et al. (2024), and others, can be understood within our framework as suboptimal nonlinear memory as well.
@@
For many kinds of sequential data, such as time-series data and text data, recent tokens are more informative of predicting the next token. Thus a natural requirement is to encourage our memory ${m}_{t}$ to focus more on the recent history by downweighting the past. Fortunately, this can be straightforwardly enforced through classic weighted linear least squares regression together with geometrically decaying weights to forget older key-value pairs.
To develop an architecture from weighted linear least squares, we only need to add weights to the optimization objective relative to ordinary least squares:
@@
    Here, to connect to past literature, we focus on weights are geometrically decaying in time: $\gamma_{i}^{(t)}=\prod_{j=i+1}^{t}\gamma_{j}$ and $\gamma_{i}\in[0,1]$ for $i=1,\ldots,t$. Each association is down-weighted based on how far it is from the current timestep $t$: $\gamma_{1}^{(t)}\leq\ldots\gamma_{t-1}^{(t)}\leq\gamma_{t}^{(t)}=1$.
@@
We can solve weighted linear regression by instead solving an unweighted regression problem involving rescaled keys and values, $\mathbf{V}_{t}\mapsto\bm{\Gamma}_{t}^{1/2}\mathbf{V}_{t}$ and $\mathbf{K}_{t}\mapsto\bm{\Gamma}_{t}^{1/2}\mathbf{K}_{t}$, where we defined a diagonal matrix $\bm{\Gamma}_{t}\in\mathbb{R}^{t\times t}$ with entries $(\bm{\Gamma}_{t})_{ii}=\gamma_{i}^{(t)}$. Making the rescaled substitution and reusing the analytical min-norm solution that we had before, we see that
$\mathbf{M}_{t}$ =$=\begin{cases}\mathbf{V}_{t}{}^{\top}\bm{\Gamma}_{t}\mathbf{K}_{t}(\mathbf{K}_{t}{}^{\top}\bm{\Gamma}_{t}\mathbf{K}_{t})^{-1},\ \text{overdetermined, e.g. }\,t\geq{D_{k}}\\ \mathbf{V}_{t}{}^{\top}(\mathbf{K}_{t}\mathbf{K}_{t}{}^{\top})^{-1}\mathbf{K}_{t},\ \text{underdetermined, e.g. }\ t<{D_{k}}.\end{cases}$ (15)
The underdetermined case has the same solution in both the weighted and unweighted case; intuitively this is because when the system is underdetermined, there are infinitely many solutions that interpolate the data points and attain zero regression loss, rendering the weights irrelevant.
@@
Once again, computing $\mathbf{M}_{t}$ requires matrix inversion which is not efficient on matrix-multiplication based hardware. Making a crude approximation similar to before of $\mathbf{K}_{t}{}^{\top}\bm{\Gamma}_{t}\mathbf{K}_{t}\approx\mathbf{I}$ results in $\mathbf{M}_{t}\approx\mathbf{V}_{t}{}^{\top}\bm{\Gamma}_{t}\mathbf{K}_{t}=\sum_{i=1}^{t}\gamma_{i}^{(t)}\mathbf{v}_{i}\mathbf{k}_{i}{}^{\top}$. We can further unroll this equation into a recurrence due to the multiplicative structure of our weights:
@@
This is exactly the recurrence implemented by gated variants of linear attention, such as GLA (Yang et al., 2024b), RetNet (Sun et al., 2023), Gateloop (Katsch, 2024), RWKV-6 (Peng et al., 2024), mLSTM (Beck et al., 2024), LRU (Orvieto et al., 2023), as well as state-space models like Mamba-2 (Dao and Gu, 2024). Forgetting older tokens significantly improves performance, which was first discovered empirically (Hochreiter and Schmidhuber, 1997, Greff et al., 2017, van der Westhuizen and Lasenby, 2018). Our derivation of gating from weighted least squares provides a mathematical grounding for the forgetting heuristic upon which these sequence layers were based.
@@
A computationally cheaper alternative is to apply a first-order gradient-descent method. We consider this here with the following design choices:
@@
    Choice of weights: assign equal weight $\gamma_{i}^{(t)}=1$ to each association,
@@
$\nabla\mathcal{L}_{t}(\mathbf{M})$ $=\sum_{i=1}^{t}\nabla L_{i}(\mathbf{M})=\sum_{i=1}^{t}(\mathbf{M}\mathbf{k}_{i}-\mathbf{v}_{i})\mathbf{k}_{i}{}^{\top}=(\mathbf{M}\mathbf{K}_{t}{}^{\top}-\mathbf{V}_{t}{}^{\top})\mathbf{K}_{t}.$ (19)
to find $\mathbf{M}_{t}$. Let the initialization be at the origin $\mathbf{M}^{(0)}_{t}=\mathbf{0}$. Iteratively solving this objective with full-batch gradient descent and step size $\beta_{t}^{(i)}$ at each iteration results in
$\mathbf{M}^{(i)}_{t}$ $=\mathbf{M}^{(i-1)}_{t}-\beta_{t}^{(i)}\nabla\mathcal{L}_{t}(\mathbf{M}^{(i-1)}_{t})$ (20)
$=\mathbf{M}^{(i-1)}_{t}-\beta_{t}^{(i)}\sum_{j=1}^{t}\nabla L_{j}(\mathbf{M}^{(i-1)}_{t})$ (21)
$=\mathbf{M}^{(i-1)}_{t}-\beta_{t}^{(i)}\sum_{j=1}^{t}(\mathbf{M}^{(i-1)}_{t}\mathbf{k}_{j}-\mathbf{v}_{j})\mathbf{k}_{j}{}^{\top}$ (22)
$=\mathbf{M}^{(i-1)}_{t}(\mathbf{I}-\sum_{j=1}^{t}\beta_{t}^{(i)}\mathbf{k}_{j}\mathbf{k}_{j}{}^{\top})+\beta_{t}^{(i)}\sum_{j=1}^{t}\mathbf{v}_{j}\mathbf{k}_{j}{}^{\top}.$ (23)
@@
Linear attention and variants as one-step of batch gradient descent
Consider the $I=1$ case with step size $\beta_{t}^{(i)}=1$. Since we initialize at the origin, we obtain the equation for linear attention
$\mathbf{M}^{(1)}_{t}=\mathbf{M}^{(0)}_{t}-\nabla\mathcal{L}_{t}(\mathbf{M}^{(0)}_{t})=\mathbf{V}_{t}{}^{\top}\mathbf{K}_{t}=\sum_{i=1}^{t}\mathbf{v}_{i}\mathbf{k}_{i}{}^{\top}.$ (24)
This is exactly the result of performing one-step of full batch gradient descent on the least square objective $\mathcal{L}_{t}$. A similar result can be shown for layers that implement linear attention with forgetting, performing one step of full batch gradient descent on the weighted least square objective.
A standard technique to accelerate the convergence of gradient descent is to precondition the gradient (Boyd and Vandenberghe, 2004). The optimal preconditioner in this case, due to the quadratic objective, is the constant Hessian of $\mathcal{L}_{t}$, $\mathbf{H}_{\mathcal{L}_{t}}(\mathbf{M})=\mathbf{K}_{t}{}^{\top}\mathbf{K}_{t}$. This preconditioner allows gradient descent to converge in a single update
$\mathbf{M}^{(1)}_{t}=\mathbf{M}^{(0)}_{t}-\nabla\mathcal{L}_{t}(\mathbf{M}^{(0)}_{t})\mathbf{H}_{\mathcal{L}_{t}}(\mathbf{M}^{(0)}_{t})^{-1}=\mathbf{V}_{t}{}^{\top}\mathbf{K}_{t}(\mathbf{K}_{t}{}^{\top}\mathbf{K}_{t})^{-1},$ (25)
@@
By taking this iterative optimization perspective, we see that linear attention and the ideal linear associative memory lie on two extremes: one-step of batch gradient descent with no preconditioning (Equation 2) versus one-step of batch gradient descent with perfect preconditioning (Equation 11). The difference in performance comes from whether the optimizer accounts for the curvature of the objective, governed by the covariance of the keys (Boyd and Vandenberghe, 2004). Exploring other preconditioners, e.g. quasi-Newton methods, that interpolates between the two extremes may be an interesting direction for future research.
Stochastic gradient descent as a layer.
@@
An alternative to full-batch gradient descent is stochastic gradient descent (SGD) where we use a subset of the samples to compute the gradient at each step. This is particularly convenient in causal sequence modeling where keys and values typically are presented one at a time, rather than all at once. In this streaming setting, we can instead perform single-example SGD to iteratively minimize Equation 8. In the case of linear memories ${\mathcal{M}_{\textsc{linear}}}$, Equation 21 with a single-example and variable step sizes $\beta_{t}^{(i)}$ becomes
$\mathbf{M}_{t}^{(i)}$ $=\mathbf{M}_{t}^{(i-1)}-\beta_{t}^{(i)}\nabla L_{t}(\mathbf{M}_{t}^{(i-1)})$ (26)
$=\mathbf{M}_{t}^{(i-1)}+\beta_{t}^{(i)}(\mathbf{v}_{t}-\mathbf{M}_{t}^{(i-1)}\mathbf{k}_{t})\mathbf{k}_{t}{}^{\top}$ (27)
$=\mathbf{M}_{t}^{(i-1)}(\mathbf{I}-\beta_{t}^{(i)}\mathbf{k}_{t}\mathbf{k}_{t}{}^{\top})+\beta_{t}^{(i)}\mathbf{v}_{t}\mathbf{k}_{t}{}^{\top},$ (28)
where $\mathbf{M}_{t}$ is the matrix representation of ${m}_{t}$. When we take only one gradient update per timestep $t$ and initialize our iterate using the previous timestep’s linear map, $\mathbf{M}_{t}^{(0)}=\mathbf{M}_{t-1}$, we end up with the recurrence of DeltaNet (Schlag et al., 2021, Yang et al., 2024c) from Equation 6:
$\mathbf{M}_{t}=\mathbf{M}_{t}^{(1)}=\mathbf{M}_{t-1}(\mathbf{I}-\beta_{t}\mathbf{k}_{t}\mathbf{k}_{t}{}^{\top})+\beta_{t}\mathbf{v}_{t}\mathbf{k}_{t}{}^{\top}.$ (29)